On 1 July 2019, BBA (pvt) Ltd borrowed Rs. 5 Million for one year, for finance the construction of a new Plant & Machinery. The interest rate on the loan is 6% and is payable on maturity of the loan. Construction work started immediately from 1 July 2019. The asset was available for use on 30 June 2020 having a construction cost of Rs. 6 Million. What is the carrying amount of the Plant & Machinery in BBA (pvt) Ltd's statement of financial position as at 30 June 2020?

 Rs. 6 Million

Rs. 6.1 Million

Rs. 6.2 Million

Rs. 6.23 Million        


1
Expert's answer
2021-06-22T15:24:03-0400

Carrying Amount is the amount which an asset is recognized after deducting any accumulated depreciation or loss.

Amount borrowed = Rs 5 million.

Amount not borrowed = Rs (6M - 5M)

=Rs 1 million.

Total interest = 6/100% × Rs 5 million

= 0.3 Million

Acquisition cost =

Rs (5 + 1 + 0.3) million = Rs 6.3 million.

Carrying amount = Rs (6.3 - 0.3) million

= Rs 6 Million.